Show that R and C and, more generally, R^n and C^n are locally compact.


1
Expert's answer
2021-07-16T09:09:47-0400

First of all, let us remark that for all "n\in \mathbb{N}", "\mathbb{C}^n" has the topology of "\mathbb{R}^{2n}" (with the canonical homeomorphism "(z_1,...,z_n)\to (\Re z_1, \Im z_1, ..., \Re z_n, \Im z_n)"), so we can just study the case of "\mathbb{R}^n".

Let "x\in \mathbb{R}^n", we consider an open neighbourhood "x\in B(x,1)" (an open ball of radius 1) and we shall prove that the unit cube "C_x =\{y=(y_1,...,y_n)\in \mathbb{R}^n \;|\; \max_{1\leq i \leq n} (|y_i-x_i| ) \leq 1\}" is compact. First we will remark that "B(x,1)\subseteq C_x", as "y\in B(x,1) \Rightarrow \sum (y_i-x_i)^2<1". Secondly we remark that a cube "C_x" is a product of "n" closed intervals "C_x=\prod_{1\leq i\leq n} [x_i-1; x_i+1]". Closed bounded intervals are compact (by Bolzano-Weierstrass theorem and the equivalence between sequential compactness and compactness for metric spaces). From this we conclude that "C_x" is compact by using one of the following arguments :

  • by Tychonoff's theorem, product of compact spaces is compact,
  • as product is finite, we can apply a weaker result - we can just use Bolzano-Weierstrass theorem for each coordinate,
  • we could have used a general result that closed bounded subsets of "\mathbb{R}^n" are compact.

Therefore, "x" admits a compact neighbourhood and thus for all "n\in \mathbb{N}", "\mathbb{R}^n" (and therefore "\mathbb{C}^n") are locally compact.